November 1996  ICTCM IX (Reno)  8

Example: The Parachute Problem

Drucker (College Mathematics Journal, May 1995)

A parachutist exits an airplane 1200 ft above the ground with zero vertical velocity. He wishes to land with a velocity of 40 ft/sec. The skydiver is assumed to weigh 240 pounds, and the problem is given assuming drag force due to air resistance proportional to velocity. The proportionality constant is taken as 2 before the parachute opens and 10 afterward. The question is then, at what time after exiting the plane should the diver open the parachute in order to land with a velocity of 40 ft/sec?